### v2

@@ 4.4 Training Schema
To improve robustness, all methods use the same physics randomization [49] at episode reset, with parameters sampled uniformly from the ranges in Table 3. Unless otherwise noted, the randomized values remain fixed for the entire episode.
@@
For obstacle-based scenarios, density starts from an easier setting and linearly ramps to the target distribution over training iterations [50]. Curriculum scheduling is identical across methods.
@@ 5.1 Evaluation Setup
  • Transformer Proprio-Vision. Self-attention over visual tokens plus a proprio token, representing attention-based cross-modal fusion [12].
@@ 5.3 Ablation Studies
Effectiveness of the Mamba cross-modal backbone. The fusion backbones is evaluated under matched modality settings. In the proprioception–vision setting, LocoMamba outperforms Transformer Proprio-Vision on all metrics in Table 4. The return rises from $511.96\pm 247.30$ to $762.34\pm 156.53$ (a $48.9\%$ increase), the number of collisions drops from $141.83\pm 158.47$ to $72.53\pm 79.47$ (a $48.9\%$ reduction), and the progress distance increases from $24.85\pm 7.34$ to $32.41\pm 5.23$ (a $30.4\%$ increase). In the vision-only setting, both methods perform poorly relative to cross-modal policies. Mamba Vision-Only remains substantially below Transformer Vision-Only in return and distance, underscoring the importance of proprioception for egocentric perception and control. These comparisons indicate that the Mamba SSM backbone is most beneficial when fusing modalities within a single policy.
Efficiency of the Mamba cross-modal backbone. Table 5 summarizes learning-curve statistics under identical training schedules. In the proprioception–vision setting, the Mamba backbone achieves a higher final reward (737.4 compared with 714.1), a steeper early-learning slope (6.41 per epoch compared with 3.66, a $75\%$ increase), higher learning efficiency (1.44 compared with 1.28), and a larger area under the learning curve per epoch (601.7 compared with 529.8). In the vision-only setting, the Mamba backbone attains a higher final reward (676.0 compared with 577.8) and a steeper early slope (4.22 compared with 3.69), with a comparable area under the curve. Taken together, these results support RQ3: the Mamba-based backbone improves optimization dynamics under the same training budget and confers efficiency benefits that translate into faster learning and stronger final performance when cross-modal fusion is used. In addition, the learning curves in Fig. 4 show faster convergence and lower variance for the Mamba-based models, indicating more stable optimization. This further supports the efficiency advantage of the Mamba SSM backbone (RQ3) and the robustness of the proposed PPO training protocol (RQ4).
@@ 5.5 Generalization to Unseen Conditions
The policies are trained in the Thin Obstacle environment and evaluate them zero-shot on two previously unseen scenarios: a rugged terrain and a dynamic-obstacle terrain with randomly moving hazards. Tables 6 and 7 report means and standard deviations over 10 runs. For vision-only policies, collision counts are omitted because the agents move too little for this metric to be informative.
@@ 6 Conclusion
Due to current budgetary and hardware-access constraints, the real-world experiments have not been conducted. When resources permit, LocoMamba will be deployed on a quadruped platform and assess sim-to-real transfer, latency, and safety under field conditions.

### v3

@@ 4.4 Training Schema
To improve robustness, all methods use the same physics randomization [mehta2020active] at episode reset, with parameters sampled uniformly from the ranges in Table 3. Unless otherwise noted, the randomized values remain fixed for the entire episode.
@@
For obstacle-based scenarios, density starts from an easier setting and linearly ramps to the target distribution over training iterations [wang2021survey]. Curriculum scheduling is identical across methods. In our experiments, the obstacle density was gradually increased in a linear schedule from a sparse to a fully dense configuration, and the exact progression was determined empirically to maintain training stability. The purpose of this design is to introduce environmental complexity smoothly rather than abruptly, ensuring that the agent can progressively adapt to denser obstacle distributions.
@@ 5.1 Evaluation Setup
  • Transformer Proprio-Vision. Self-attention over visual tokens plus a proprio token, representing attention-based cross-modal fusion [yang2021learning].
@@ 5.3 Ablation Studies
Effectiveness of the Mamba cross-modal backbone. The fusion backbones is evaluated under matched modality settings. In the proprioception–vision setting, LocoMamba outperforms Transformer Proprio-Vision on all metrics in Table 4. The return rises from $511.96\pm 247.30$ to $762.34\pm 156.53$ (a $48.9\%$ increase), the number of collisions drops from $141.83\pm 158.47$ to $72.53\pm 79.47$ (a $48.9\%$ reduction), and the progress distance increases from $24.85\pm 7.34$ to $32.41\pm 5.23$ (a $30.4\%$ increase). In the vision-only setting, both methods perform poorly relative to cross-modal policies. Mamba Vision-Only remains substantially below Transformer Vision-Only in return and distance, underscoring the importance of proprioception for egocentric perception and control. The lower performance of the Mamba Vision-Only model arises from its higher temporal sensitivity: without proprioceptive input, its recurrent state gradually drifts under partial observability. In contrast, the Transformer’s self-attention captures limited historical context within each visual window. This highlights that Mamba’s selective recurrence benefits most from proprioceptive grounding rather than indicating a structural weakness. These comparisons indicate that the Mamba SSM backbone is most beneficial when fusing modalities within a single policy.
Efficiency of the Mamba cross-modal backbone. Table 5 summarizes learning-curve statistics under identical training schedules. In the proprioception–vision setting, the Mamba backbone achieves a higher final reward (737.4 compared with 714.1), a steeper early-learning slope (6.41 per epoch compared with 3.66, a $75\%$ increase), higher learning efficiency (1.44 compared with 1.28), and a larger area under the learning curve per epoch (601.7 compared with 529.8). In the vision-only setting, the Mamba backbone attains a higher final reward (676.0 compared with 577.8) and a steeper early slope (4.22 compared with 3.69), with a comparable area under the curve. Taken together, these results support RQ3: the Mamba-based backbone improves optimization dynamics under the same training budget and confers efficiency benefits that translate into faster learning and stronger final performance when cross-modal fusion is used. In addition, the learning curves in Fig. 4 show faster convergence and lower variance for the Mamba-based models, indicating more stable optimization. As shown in Table 6, the proposed LocoMamba model achieves the highest throughput (104.7 samples/s) while maintaining a lower allocated memory footprint (207 MB) compared to its Transformer-based counterpart (226 MB). Moreover, despite having slightly fewer parameters (300K vs. 353K), LocoMamba exhibits comparable FLOPs per step, demonstrating that its superior runtime efficiency results from architectural optimization rather than reduced computational load. This further supports the efficiency advantage of the Mamba SSM backbone (RQ3) and the robustness of the proposed PPO training protocol (RQ4).
@@
Table 6: Computational and Memory Efficiency Comparison between Mamba-based and Transformer-based Models
Model Reserved Memory (MB) Allocated Memory (MB) Samples/s FLOPs/step (Policy) FLOPs/step (Value) Params (Policy) Params (Value)
LocoMamba (Ours) 3160 207.2 104.68 4113664 4112384 300070 298785
Transformer Proprio-Vision 3156 226.4 98.87 5012992 5011712 353190 351905
Mamba Vision-Only 3098 189.6 100.35 4198912 4197632 193254 191969
Transformer Vision-Only 3022 199.2 102.84 4826624 4825344 232806 231521
To further quantify the trade-off between spatial resolution and computational efficiency, we conducted an ablation study by varying the number of visual tokens ($32$, $64$, and $256$) generated from the CNN patchification process. The corresponding performance and efficiency results are reported in Tables 7 and 8, respectively. As expected, increasing the token count provides the model with finer spatial granularity, which improves perception accuracy in trained environments. However, it also leads to higher computational overhead due to the quadratic growth of feature representations and the additional intermediate states required in the Mamba layers. Among the tested settings, the $256$-token configuration yields the highest return and distance on the trained (thin obstacle) terrain but exhibits greater memory usage and degraded generalization on unseen rugged terrains. In contrast, the $64$-token configuration achieves a favorable balance, offering competitive performance across all environments while maintaining moderate GPU memory consumption and FLOPs. These results indicate that an intermediate token size provides the best trade-off between spatial fidelity, computational efficiency, and cross-domain robustness.
Table 7: Performance comparison across different visual token configurations (32, 64, 256).
Token Size Reward Collisions Distance
Thin Obstacle Environment
32 740.55 $\pm$ 193.45 57.88 $\pm$ 39.77 32.93 $\pm$ 7.39
64 (Ours) 762.34 $\pm$ 156.53 72.53 $\pm$ 79.47 32.41 $\pm$ 5.23
256 797.46 $\pm$ 111.29 106.95 $\pm$ 92.71 34.74 $\pm$ 5.02
Dynamic Obstacle Environment
32 697.33 $\pm$ 587.19 12.31 $\pm$ 89.88 29.39 $\pm$ 8.00
64 (Ours) 749.09 $\pm$ 269.19 38.47 $\pm$ 43.96 32.51 $\pm$ 5.84
256 682.21 $\pm$ 569.84 260.51 $\pm$ 203.64 29.08 $\pm$ 9.81
Rugged Unseen Environment
32 563.39 $\pm$ 166.65 1842.38 $\pm$ 357.57 23.97 $\pm$ 5.78
64 (Ours) 583.76 $\pm$ 154.57 470.27 $\pm$ 108.15 24.04 $\pm$ 6.18
256 562.61 $\pm$ 185.64 4166.64 $\pm$ 708.69 23.59 $\pm$ 6.70
Table 8: Efficiency comparison across different visual token configurations (32, 64, 256).
Token Size Memory Usage (Reserved / Allocated) FLOPs (Policy / Value) Params (Policy / Value)
32 2980 MB / 200 MB 3,901,952 / 3,903,232 263,137 / 264,422
64 3160 MB / 207.2 MB 4,112,384 / 4,113,664 298,785 / 300,070
256 3546 MB / 359.6 MB 6,928,640 / 6,929,920 604,065 / 605,350
@@ 5.5 Generalization to Unseen Conditions
The policies are trained in the Thin Obstacle environment and evaluate them zero-shot on two previously unseen scenarios: a rugged terrain and a dynamic-obstacle terrain with randomly moving hazards. Tables 9 and 10 report means and standard deviations over 10 runs. For vision-only policies, collision counts are omitted because the agents move too little for this metric to be informative.
@@ 6 Conclusion
Due to current budgetary and hardware-access constraints, the real-world experiments have not been conducted. When resources permit, LocoMamba will be deployed on a quadruped platform and assess sim-to-real transfer, latency, and safety under field conditions. Future work will incorporate more realistic visual perturbations such as occlusion holes, depth bias, and motion blur, as well as temporal latency and misalignment, to further evaluate the robustness of the Mamba-based cross-modal fusion in real-world conditions.
@@
Acknowledgment
This research was funded by the Graduate Innovation Fund of Jilin University under Grant 101832020CX130.